A student receives his grade report from a local community college, but the GPA is smudged. He took the following classes: a 2 hour credit art, a 3 hour credit history, a 4 hour credit science course, a 3 hour credit mathematics course, and a 1 hour science lab. He received a “B” in the art class, an “A” in the history class, a “C” in the science class, a “B” in the mathematics class, and an “A” in the science lab. What was his GPA if the letter grades are based on a 4 point scale? (A=4, B=3, C=2, D=1, F=0). Please help.
Answer by bucky(2100)

This is a problem involving "weighted" grades. You solve it by multiplying the credit for
a course times the numerical value of the grade received, then adding up all the credits
and dividing it into the sum of all the grade products.
.
Hard to explain, but easy to do.
.
Art .......... 2 credit hours times B (3 points) = 6 grade points
History ...... 3 credit hours times A (4 points) = 12 grade points
Science ...... 4 credit hours times C (2 points) = 8 grade points
Mathematics .. 3 credit hours times B (3 points) = 9 grade points
Science Lab .. 1 credit hour times A (4 points) = 4 grade points
.
Now add up all the credit hours (2 + 3 + 4 + 3 + 1 = 13 hours)
.
Then add up all the grade points earned (6 + 12 + 8 + 9 + 4 = 39 grade points)
.
Finally divide the credit hours into the grade points and you have the grade point average (GPA)
.
GPA = 39/13 = 3.00
.
The grade point average (GPA) of this student is 3 point oh oh ... 3.00.
.
Hope this makes the computation of grade point average understandable. One thing it should
make clear to you is that it is important that you get a good grade in courses that have
the higher credit hours values. An A in a course that has a credit hour rating of 4 or 5
scores you more grade points than an A in a 1 credit hour course and therefore an A
in the 4 or 5 hour course really helps to increase your GPA.